The 8-bit registers AR, BR, CR, and DR initially have the following values: AR= 11110010, BR= 11101111, CR= 10111001 , DR= 11101010 Determine the 8-bit values in each register after the execution of the following sequence of microoperations. AR <-AR + BR CR <- CR 16 ^ DR, BR <- BR + 1 AR <-AR – CR


1
Expert's answer
2021-10-04T02:26:50-0400

"AR<-(-AR)+BR\\
-AR=00001101+1=00001110\\
BR=11101111\\
AR<--AR+BR=11111101\\
CR<-(-CR)=01000111\\
DR<-16{}\hat{}DR=00010000AND11110010=00010000\\
BR<-BR+1=00010001+1=00010010\\
AR<-(-AR)+(-CR)=00000010+10111001=10111011"